OpenStudy (anonymous):

I am having trouble with this physics question, can someone please help me how to solve this question? If a spacecraft is 18,927,000,000 km from the Sun and is traveling at 1,472,400 km/day........ What is the spacecraft's distance from the Sun in one (1) day? -and- What is the spacecraft's distance from the Sun in the second (2) day?

OpenStudy (anonymous):

Is there a direction given along with the velocity?

OpenStudy (anonymous):



OpenStudy (anonymous):

after one day= V* day= 1,472,400 (km/day) * 1 (day)= 1,472,400 km distance from the sun after one day = 18,927,000,000-1,472,400= distance from the sun after two days = 18,927,000,000- 2*1,472,400=

OpenStudy (mrnood):

OMG - do the people who set these questions even LOOK at what they said? This question is SO poorly worded - it just serves to totally confuse the student. As a PHYSICS question it fails in 3 MAJOR points: 1) it doesn't tell you the direction of the velocity. It could be travelling directly towards OR away from the sun - or anything in between. Indeed it could be travelling tangentially to the sun in an orbit (I haven't worked out the necessary speed to be in orbit at this radius - but in any case you CANNOT answer tis question without making a totally unjustified assumption. 2) A body in such a position is 126 AU from the sun - the Solar system is approx. 90 AU radius, so this 'spaceship' has left the solar system. Even the Voyager craft which were launched 40 years ago are only JUST at the edge of the solar system. Maybe they are training you to look for little green men. 3) Whilst this is essentially 'interstellar' space the craft is likely to be accelerating. It does not tell you that the speed is constant. I suggest you take this question back to class and ask them for a question based on the rigour of Physics.